Random samples of size 36 are taken from a large population whose mean is 120 and standard deviation is 39. The mean and standard error of the sampling distribution ofsample mean, respectively, are:
Answer #1
mean error = 120/36 = 3.333

standard deviation error =39/36 =1.08
answered by: Kathryn
Answer #2

mean=120

standard error= s/√n

=39/6

=6.5

answered by: Fatih